OpenStudy (anonymous):

>> WILL FAN AND GIVE MEDAL << The graph of one particular link includes the two points (5,8) and (10,6). Find the slope of the line. A. 2/5 B. -2/5 C. 5/2 D.-5/2

OpenStudy (anonymous):

@MathsPro @amistre64

OpenStudy (fibonaccichick666):

what is the formula for slope?

OpenStudy (anonymous):

@FibonacciChick666 Label the points as x1 = 1, y1 = - 4, x2 = -4, and y2 = 2. To find the slope m of the line segment joining the points, use the slope formula :

OpenStudy (fibonaccichick666):

what is the slope formula is my question

OpenStudy (anonymous):

@FibonacciChick666 Look it up.

OpenStudy (fibonaccichick666):

no. I am asking you because YOU need to know it. I already do.

OpenStudy (fibonaccichick666):

and please stop tagging me each time, it's not necessary

OpenStudy (anonymous):

@FibonacciChick666 I already know it, I want help, not an attitude.

OpenStudy (fibonaccichick666):

then tell me it. like I asked

OpenStudy (anonymous):

I already know HOW to do it. I just got the WRONG ANSWER. I got, -3/4

OpenStudy (fibonaccichick666):

then tell me how you did it, so I can find where you went wrong

OpenStudy (anonymous):

Okay, so first I did. x1, y1 x1, y1 Then I plugged in the numbers. m= 5-8= -3 10-6= 4 -3/4

OpenStudy (fibonaccichick666):

ok, so first question, can you tell me what you used for x1, y1,x2,y2?

OpenStudy (fibonaccichick666):

It didn't copy

OpenStudy (anonymous):

I used: (5,8), (10,6)

OpenStudy (fibonaccichick666):

can you type it like so: x1=___ y1=____ x2=____ y2=____

OpenStudy (anonymous):

x1= 5 y1= 8 x2= 10 y2= 6

OpenStudy (fibonaccichick666):

k so now, you did not apply the formula correctly, so that is why I am going to ask for the slope formula again

OpenStudy (fibonaccichick666):

So, \[m=\frac{y_1-y_2}{x_1-x_2}\] so can you input your values directly into this^^ and tell me that?/

OpenStudy (fibonaccichick666):

just right click copy and paste the latex commands

OpenStudy (anonymous):

I don't know how.

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

you don't know how to do the plug and chug or you don't know how to type it all pretty like here?

OpenStudy (fibonaccichick666):

because if it's just the typing, you can draw it out too

OpenStudy (anonymous):



OpenStudy (fibonaccichick666):

ah, that is the problem, look at what you put in, You did \[m=\frac{x_1-y_1}{x_2-y_2}\] but you want \[m=\frac{y_1-y_2}{x_1-x_2}\]